OpenStudy (anonymous):

Simplify the expression below. -5-8

OpenStudy (anonymous):

-5 + (-8) = ....

OpenStudy (anonymous):

what did you do

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

shoot never mind lol

OpenStudy (anonymous):

-5x-8

OpenStudy (anonymous):

subtracting is the same as adding the opposite. so -5 - (8) = -5 + (-8) = -13

OpenStudy (anonymous):

okay so i was right, got confused for a second.

OpenStudy (anonymous):

lol i see and thank you.. how about √ 169

OpenStudy (anonymous):

13

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

typed it into my scientific calculator :)

OpenStudy (anonymous):

(; ok